The following transactions and adjusting entries were completed by a paper-packaging company called Gravure Graphics International during 2018 and 2019. The company uses straight-line depreciation for trucks and other vehicles, double-declining-balance depreciation for buildings, and straight-line amortization for patents.

2018
January 2 Paid $94,000 cash to purchase storage shed components.
January 3 Paid $4,000 cash to have the storage shed erected. The storage shed has an estimated life of 10 years and a residual value of $7,000.
April 1 Paid $37,000 cash to purchase a pickup truck for use in the business. The truck has an estimated useful life of five years and a residual value of $5,000.
May 13 Paid $300 cash for minor repairs to the pickup truck's upholstery.
July 1 Paid $21,000 cash to purchase patent rights on a new paper bag manufacturing process. The patent is estimated to have a remaining useful life of five years.
December 31 Recorded depreciation and amortization on the pickup truck, storage shed, and patent.
2019
June 30 Sold the pickup truck for $31,000 cash. (Record the depreciation on the truck prior to recording its disposal.)
December 31 Recorded depreciation on the storage shed. Also determined that the patent was impaired and wrote off its remaining book value (i.e., wrote down the book value to zero).

Required:

Prepare the journal entries required on each of the above dates. (If no entry is required for a transaction/event, select "No Journal Entry Required" in the first account field. Do not round intermediate calculations.)
Answer #1
No. Date General Journal Debit Credit
1 January 02, 2018 Buildings 94000
Cash 94000
(Storage shed components purchased for cash)
2 January 03, 2018 Buildings 4000
Cash 4000
(Erection costs paid in cash)
3 April 01, 2018 Equipment 37000
Cash 37000
(Pickup truck purchased for cash)
4 May 13, 2018 Repairs and maintenance expense 300
Cash 300
(Repairs to pickup truck paid in cash)
5 July 01, 2018 Patent 21000
Cash 21000
(Purchase of patents for cash)
6 December 31, 2018 Depreciation expense 24400
Amortization expense ($21000/5 years x 6/12) 2100
Accumulated depreciation-Building 19600
Accumulated depreciation-Equipment 4800
Accumulated amortization-Patent 2100
(Depreciation and amortization expense for the year)
7 June 30, 2019 Depreciation expense 3200
Accumulated depreciation-Equipment 3200
(Depreciation expense on truck)
8 June 30, 2019 Cash 31000
Accumulated depreciation-Equipment 8000
Equipment 37000
Gain on disposal 2000
(Sale of truck for cash)
9 December 31, 2019 Depreciation expense 15680
Accumulated depreciation-Building 15680
(Depreciation expense on building)
10 December 31, 2019 Impairment loss 18900
Patent 18900
(Loss on impairment of patent)

Workings:

Straight-line depreciation on truck:

Annual depreciation = (Cost - Salvage value)/Estimated useful life = ($37000 - $5000)/5 years = $32000/5 = $6400

2018 depreciation for 9 months = $6400 x 9/12 = $4800

2019 depreciation for 6 months = $6400 x 6/12 = $3200

Double-declining depreciation on storage shed (Building):

Straight-line annual depreciation rate = 100%/Estimated useful life = 100%/10 years = 10%

DDB rate = 2 x 10% = 20%

Year Book Value Beginning of Year x Depreciation Rate = Annual Depreciation Expense End of Year
Accumulated Depreciation Book Value
2018 98000 20% 19600 19600 78400
2019 78400 20% 15680 35280 62720